I need help making a table and putting the right numbers in the right spots. I also have to know how to put an equation together to solve. Here is a Story problem.
How many pounds of candy selling at $3.75/lb should be mixed with 35 lbs of candy selling at $5/lb to produce a mixture that should sell at $4.25/lb?
Here is the table I made.

lbs $/lb total $ for candy
Candy 1 x 3.75 $3.75x
Candy2 35 5.00 35(5)=175
mix 35+x 4.25 (35+x)(4.25)


Can you please tell me what I did wrong and show me where the right numbers go in the right spots? Please show me the equation too?
Thanks

Answer by scott8148(6628)

your table looks good

the equation is in the last column ___ ($3.75)x + 175 = (35 + x)($4.25)

3.75x + 175 = 148.75 + 4.25x

26.25 = .5x

52.5 = x